OpenStudy (anonymous):

its 56 1/4at a percent%... so whats the fraction

jimthompson5910 (jim_thompson5910):

56 1/4 = 56 + 1/4 = 224/4 + 1/4 = 225/4 So 56 1/4 = 225/4 So 56 1/4% = 225/4% Now divide the right side by 100 to get 225/400 and reduce that to get 9/16 So 56 1/4% as a fraction is 9/16

OpenStudy (anonymous):

\[56\text{ }1/4\text{ %}=56.25\text{ %}=\frac{5625}{100}=\frac{225}{4} \]